Solve.
Mark drove 250 miles to visit Sandra. Returning by a shorten route, he found that the trip was only 225 miles, but traffic slowed his speed by 5 miles. If the two trips took exactly the same time, what was his rate on the return trip?

Answer by checkley71(8403)

250/R=225/(R-5) NOW CROSS MULTIPLY
250(R-5)=225R
250R-1250=225R
250R-225R=1250
25R=1250
R=1250/25
R=50 MPH IS THE SPEED OF THE FIRST TRIP
50-5=45 MPH FOR THE RETURN TRIP.
PROOF
250/50=225/45
5=5